OpenStudy (anonymous):

Steven recorded the growth of a plant over 10 wk for his science project. He made a graph that shows how much the plant grew each week. The plant was 2 in. tall when Steven started his project. At week 5, the plant was 8 in. tall.

OpenStudy (anonymous):

Complete the equation that represents Steven’s situation. y = ? x + ?

OpenStudy (butterflydreamer):

firstly, can you find the slope of the line? :)

OpenStudy (anonymous):

6/5

OpenStudy (anonymous):

i'm thinking this is the answer y = 6/5 x +1.2

OpenStudy (anonymous):

oops sorry i mean y = 1.2x +2

OpenStudy (anonymous):

i think

OpenStudy (butterflydreamer):

slope = 8 - 2 / 5 - 0 = 6/5 = 1.2 (basically meaning that m = 1.2) y-intercept: b = 2 y = mx + b so, y = 1.2x + 2 you're correct :) good job and sorry for my slow reply

OpenStudy (anonymous):

Thank You! :)

OpenStudy (butterflydreamer):

you're welcome :)!!